Medical aid board wants to find if the mean time to settle medical claims of members differs between 2 major medical schemes. The sample mean time to settlement was 10.8 days. Population standard deviation of 3.2 days for 1 and the other 12.4 days with population deviation 2.3 days



Formulate the null and alternative hypothesis and test this hypothesis at the 1% significance level and draw a conclusion?

1
Expert's answer
2022-03-02T11:02:29-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2"

"H_1:\mu_1\not=\mu_2"

This corresponds to a two-tailed test, and a z-test for two means, with known population standard deviations will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is "z_c=2.5758."

The rejection region for this two-tailed test is "R = \{z: |z| > 2.5758\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{X}_1-\bar{X}_2}{\sqrt{\sigma^2_1/n_1+\sigma^2_2/n_2}}"

"=\dfrac{10.8-12.4}{\sqrt{3.2^2/14+2.3^2/15}}\approx-1.5367"

Since it is observed that "|z| = 1.5367 \le 2.5758=z_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-1.5367)=0.124369," and since "p = 0.124369 \ge 0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the "\alpha = 0.01" significance level.